OpenStudy (anonymous):

what is the sum of the reciprocals of the solutions of x^3-3x^2-13x+15=0?

OpenStudy (blockcolder):

If \(r_1, r_2, r_3\) are the roots, then \[\frac{1}{r_1}+\frac{1}{r_2}+\frac{1}{r_3}=\frac{r_2r_3+r_1r_3+r_1r_2}{r_1r_2r_3}\] By Viete's formulas, \(r_2r_3+r_1r_3+r_1r_2=-13\) and \(r_1r_2r_3=15\).

OpenStudy (blockcolder):

Oops. \(r_1r_2r_3=-15\), not 15.

OpenStudy (radar):

Solve for the roots. since the factors of 15 are 3, 5 or 15, 1. The most likely candidate is the 3,5 using synthetic division or long division you will find that (x+3) is one of the factors, dividing x^3-3x^2-13x+5 by (x+3) gives you x^2-6x+5 which factors to (x-5)(x-1) we have now the three factors (x+3), (x-5)(x-1) the roots then are: x=-3, x=5, x=1 giving you reciprocals of -1/3, 1, 1/5 summing them: -1/3 +1/5 + 1= -5/15 + 3/15 + 15/15 =13/15